OpenStudy (anonymous):

tickets for the school play were $6.00 and 1/5 of each ticket paid for the cost of printing the programs. how much of each ticket went towards printing the programs

OpenStudy (anonymous):

anyone help me please

OpenStudy (anonymous):

1/5*6 = 6/5 = $1.20